# ReGATE: Learning Faster and Better with Fewer Tokens in MLLMs

### v1

@@ 4.2 Evaluation Benchmarks
| Open-source Models | |||||||
| Video-LLaVA Lin et al. (2024a) | Vicuna-7B | 8 | – | 39.9 | 39.1 | 47.3 | 38.4 |
| LLaMA-VID Li et al. (2024d) | Llama-2-7B | 1fps | – | 25.9 | – | 33.2 | 38.5 |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Vicuna-7B | 32 | – | – | 43.5 | – | 43.9 |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Qwen2-32B | 32 | – | 60.2 | – | 65.5 | 60.9 |
| VILA1.5 Lin et al. (2024b) | Llama-2-40B | 8 | – | 60.1 | – | 56.7 | 58.0 |
| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | 32 | – | 58.2 | 56.4 | 64.7 | 60.1 |
| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | – | 65.1 | 56.0 | 70.2 | 65.0 |
| VideoLLaMA3 Zhang et al. (2025) | Qwen2.5-7B | 1fps | – | 66.2 | 59.8 | 73.0 | 63.3 |
| Proprietary Models | |||||||
| Gemini-1.5-Flash Gemini et al. (2024) | – | – | – | 70.3 | 61.6 | – | 65.7 |
| Gemini-1.5-Pro Gemini et al. (2024) | – | – | – | 75.0 | 64.0 | – | 71.2 |
| GPT-4o Hurst et al. (2024) | – | – | – | 71.9 | 66.7 | 64.6 | 72.2 |
| Models w/wo ReGATE | |||||||
@@
| VideoLLaMA2-ReGATE | Qwen2-7B | 16 | 49.27M (↓ 41.22%) | 54.5 | 47.6 | 54.5 | 56.4 |
| Model | LLM | Frames | Tokens | MVBench | Perception | NeXT-QA | |
| Open-source Models | |||||||
| Video-LLaVA Lin et al. (2024a) | Vicuna-7B | 8 | – | 41.0 | 44.3 | 25.8 | – |
| LLaMA-VID Li et al. (2024d) | Llama-2-7B | 1fps | – | 41.9 | 44.6 | – | – |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Vicuna-7B | 32 | – | 46.5 | 48.8 | 25.6 | – |
| LLaVA-NeXT-Video Zhang et al. (2024b) | Qwen2-32B | 32 | – | – | 59.4 | – | 77.3 |
| VILA1.5 Lin et al. (2024b) | Llama-2-40B | 8 | – | – | 54.0 | – | 67.9 |
| LLaVA-OneVision Li et al. (2024a) | Qwen2-7B | 32 | – | 56.7 | 57.1 | 29.4 | 79.4 |
| Qwen2.5-VL Bai et al. (2025) | Qwen2.5-7B | – | – | 69.6 | 70.5 | – | – |
| VideoLLaMA3 Zhang et al. (2025) | Qwen2.5-7B | 1fps | – | 69.7 | 72.8 | – | 84.5 |
| Proprietary Models | |||||||
| Gemini-1.5-Pro Gemini et al. (2024) | – | – | – | 60.5 | – | 22.6 | – |
| GPT-4o Hurst et al. (2024) | – | – | – | 64.6 | – | 38.2 | – |
| Models w/wo ReGATE | |||||||
| VideoChat2 | Mistral-7B | 16 | 3.93B | 55.7 | 48.4 | 22.0 | 75.2 |
| VideoChat2-ReGATE | Mistral-7B | 16 | 2.22B (↓ 43.51%) | 56.6 | 50.0 | 22.8 | 75.5 |
| VideoLLaMA2 | Qwen2-7B | 16 | 83.82M | 52.0 | 53.0 | 24.6 | 70.8 |
| VideoLLaMA2-ReGATE | Qwen2-7B | 16 | 49.27M (↓ 41.22%) | 53.6 | 54.1 | 25.2 | 70.0 |
4.3 Baseline Models
We evaluate ReGATE against a comprehensive set of baselines, including the adapted VideoLLaMA2 Cheng et al. (2024) and VideoChat2 Li et al. (2024c) models. Our comparison covers a broad range of state-of-the-art open-source models, primarily drawn from high-performing families such as LLaVA and Qwen. We also report results from proprietary models in the Google Gemini, OpenAI GPT, and Anthropic Claude series. This diverse set of baselines spans multiple LLM backbones and model sizes, ensuring a robust and meaningful comparison. The specific models evaluated across image and video tasks are listed in Tables 1, 2, and 3.
@@ 4.4 Results
Learning better: ReGATE’s accuracy gains across image and video benchmarks.
@@
The comprehensive results presented in Tables 1, 2, and 3 show how VideoLLaMA2 and VideoChat2 perform, with and without ReGATE, across a range of image, short video, and long video understanding benchmarks. ReGATE improves performance consistently by focusing computation on the most informative tokens. For example, VideoLLaMA2-ReGATE outperforms the baseline VideoLLaMA2 on most tasks while using 41.22% fewer tokens. Similarly, VideoChat2-ReGATE achieves better results than the baseline VideoChat2 while using 43.51% fewer tokens.
@@
On image understanding tasks that require multimodal reasoning, both models show significant gains. VideoLLaMA2-ReGATE improves by 19.1% on ScienceQA and by up to 33.1 points on MME. VideoChat2-ReGATE improves by 5.8% and 46.1 points on the same benchmarks. For long video understanding, VideoChat2-ReGATE shows strong improvements of 6.7% on VideoMME and 4.5% on MLVU. VideoLLaMA2-ReGATE also improves, though more modestly, with gains of 0.8% and 1.3% on the same tasks. Short video tasks benefit as well. VideoLLaMA2-ReGATE improves by 1.6% on MVBench and 1.1% on Perception, while VideoChat2-ReGATE gains 0.9% and 1.6%, respectively.
Overall, these results demonstrate ReGATE’s ability to adapt across diverse tasks by efficiently directing computational resources to the most important visual and semantic content.
| Model | Tokens | Train Time | Acc. (%) |
| VideoLLaMA2 | 83.82M | 32.4h | 52.0 |
| VideoLLaMA2-ReGATE | 49.27M | 26.9h | 53.6 |
| VideoLLaMA2-ReGATE | 29.32M | 16.0h | 51.9 |
| VideoChat2 | 3.93B | 37.2h | 55.7 |
| VideoChat2-ReGATE | 2.22B | 32.5h | 56.6 |
| VideoChat2-ReGATE | 1.51B | 21.6h | 55.5 |
Learning faster: ReGATE’s efficiency gains.
Table 4 presents detailed efficiency gains in token usage, training time, and accuracy on the MVBench benchmark.
For VideoLLaMA2, ReGATE closely matches the baseline accuracy (51.9% vs. 52%) in just 16.0 hours, which is less than half the time required for standard fine-tuning (32.4 hours). It does so using only 29.32 million tokens, approximately 35% of the 83.82 million tokens used by the baseline. When training is extended to 26.9 hours (still 5.5 hours less than the baseline), ReGATE processes 41.51% fewer tokens and achieves a higher accuracy of 53.6%.
For VideoChat2, which uses parameter-efficient LoRA fine-tuning, the improvements in training time are more modest. Specifically, ReGATE closely matches the baseline accuracy (55.5% vs. 55.7%) in 21.6 hours, compared to 37.2 hours for the baseline. Furthermore, when training time increases to 32.5 hours (still 4.7 hours less than the baseline), ReGATE processes 43.51% fewer tokens (2.22 billion vs. 3.93 billion) and achieves an improved accuracy of 56.6%.
This speed-up difference between VideoLLaMA2 and VideoChat2 arises from the contrast between full and LoRA fine-tuning strategies. In full fine-tuning, as used in VideoLLaMA2, both forward and backward passes through the model are computationally expensive. By pruning tokens, ReGATE speeds up both passes, especially the backward pass where gradients are computed for all model parameters. In LoRA fine-tuning, as used in VideoChat2, most parameters are frozen, and the backward pass is already efficient since gradients are only computed for a small number of adapter parameters. While ReGATE still accelerates the forward pass through the frozen backbone, the total time savings are smaller because the backward pass is not a bottleneck. Overall, ReGATE delivers significant gains in both token efficiency and training time across different training strategies, making it a flexible and effective solution for reducing computation without compromising performance.
5 Conclusion and Future Work
We introduced ReGATE, a reference-guided token gating framework that accelerates the training of multimodal large language models. By combining a student model’s learning difficulty with reference losses from a frozen text-only teacher, ReGATE dynamically focuses computation on the most informative tokens while skipping those less relevant for multimodal understanding. The method is simple to implement, requires no architectural changes, and substantially improves training efficiency. Experiments show that ReGATE achieves comparable or better accuracy than standard full fine-tuning, using only a fraction of the tokens and significantly less training time. These gains come without compromising model quality. In fact, ReGATE consistently outperforms baselines across a wide range of image and video benchmarks, demonstrating strong data efficiency and generalization. Future work will explore adaptive scheduling for token sparsity by dynamically adjusting the retained token ratio based on task complexity, model stability, and training progress (e.g., starting with higher sparsity early and relaxing it as fine-tuning progresses). We will also investigate fine-grained sparsity control across layers or attention heads for more efficient resource allocation. Moreover, we aim to generalize the notion of “reference” beyond frozen text-only teachers. Using stronger or multimodal teacher models (e.g., vision/video-language) could provide richer supervision for gating, improving cross-modal alignment and enhancing performance on complex spatial and temporal tasks.
Limitations
Due to computational resource constraints, we validate ReGATE on 7B-parameter models, VideoLLaMA2 and VideoChat2, demonstrating clear effectiveness and efficiency gains at this scale. However, the full potential of ReGATE likely emerges with larger models (e.g., 30B or 70B+ parameters) and massive, web-scale datasets, where time and cost savings become more significant. Future work should focus on evaluating ReGATE’s performance and scalability in such high-resource settings.
Appendix
Appendix A Ablation Study
How does each signal in our scoring mechanism affect performance?
@@ Appendix A Ablation Study
To validate the contributions of the individual components within our dual-signal token scoring mechanism, we conduct an ablation study on the hyperparameter . This coefficient balances the two core signals in our difficulty score formulation: , where is the student’s dynamic EMA difficulty and is the static reference loss from the teacher model. By varying , we can isolate the impact of each signal.
We evaluate three values for : 0.0, 0.5, and 1.0. The experiments use our VideoLLaMA2-ReGATE setup with all other hyperparameters fixed for a fair comparison. As shown in Table 5, , which balances the reference loss and the student’s EMA-based difficulty, results in the best performance.
@@
| Student EMA Only | 51.3 | |
| Reference Loss Only | 51.1 | |
| Combined Signals | 53.6 |
Appendix B Qualitative Analysis of Reference Loss
To validate the core mechanism of ReGATE, we qualitatively analyze the reference loss signal that guides its token selection. We assume that a high loss score from the text-only teacher indicates that a token requires visual information to be understood. Figure 3 shows two video Q&A examples, visualizing the loss for each word in the answer as calculated by a Mistral-7B Jiang et al. (2023) teacher model.
@@ Appendix B Qualitative Analysis of Reference Loss
| SEED-Bench | Answer with the option’s letter from the given choices directly. |
@@
| VideoMME | Answer with the option’s letter from the given choices directly. |
@@
Appendix C Additional Benchmarks Details
Table 6 lists the evaluation prompts corresponding to each benchmark used in the experiments.

### v2

@@ 4.2 Benchmarks and Baselines
[…]
| Video-LLaVA [lin2024videollava] | Vicuna-7B | 8 | – | 39.9 | 39.1 | 47.3 | 38.4 |
| LLaMA-VID [li2024llamavid] | Llama-2-7B | 1fps | – | 25.9 | – | 33.2 | 38.5 |
| LLaVA-NeXT-Video [zhang2024llavanextvideo] | Vicuna-7B | 32 | – | – | 43.5 | – | 43.9 |
| LLaVA-NeXT-Video [zhang2024llavanextvideo] | Qwen2-32B | 32 | – | 60.2 | – | 65.5 | 60.9 |
| VILA1.5 [lin2024vila] | Llama-2-40B | 8 | – | 60.1 | – | 56.7 | 58.0 |
| LLaVA-OneVision [llavaov] | Qwen2-7B | 32 | – | 58.2 | 56.4 | 64.7 | 60.1 |
| Qwen2.5-VL [bai2025qwen25vl] | Qwen2.5-7B | – | – | 65.1 | 56.0 | 70.2 | 65.0 |
| VideoLLaMA3 [zhang2025videollama3] | Qwen2.5-7B | 1fps | – | 66.2 | 59.8 | 73.0 | 63.3 |
[…]
| Gemini-1.5-Flash [geminiteam2024gemini15] | – | – | – | 70.3 | 61.6 | – | 65.7 |
| Gemini-1.5-Pro [geminiteam2024gemini15] | – | – | – | 75.0 | 64.0 | – | 71.2 |
| GPT-4o [openai2024gpt4ocard] | – | – | – | 71.9 | 66.7 | 64.6 | 72.2 |
[…]
@@
[…]
| InternVL3.5 | Qwen3-14B | 16 | 3.96B | 62.4 | 57.9 | 63.7 | 64.7 |
[…]
@@ 4.3 Results
Learning better: ReGATE’s accuracy gains across image and video benchmarks. The comprehensive results presented in Tables 4, 4.2, and LABEL:tab:short-video show how VideoLLaMA2, VideoChat2, and InternVL3.5 perform, with and without ReGATE, across a range of image, short video, and long video understanding benchmarks. ReGATE improves performance consistently by focusing computation on the most informative tokens. Figure LABEL:fig:saliency visualizes the learned attention maps, showing that ReGATE concentrates attention on visually and semantically important tokens compared to standard fine-tuning. For example, VideoLLaMA2-ReGATE outperforms the baseline VideoLLaMA2 on most tasks while using 41.22% fewer tokens. Similarly, VideoChat2-ReGATE achieves better results than the baseline VideoChat2 while using 43.51% fewer tokens. At the larger 14B scale, InternVL3.5-ReGATE also shows gains across both image and video benchmarks, confirming that ReGATE remains effective for high-capacity backbones.
@@
On image understanding tasks requiring multimodal reasoning, all three models show significant gains. VideoLLaMA2-ReGATE improves by 19.1% on ScienceQA and by up to 33.1 points on MME. VideoChat2-ReGATE improves by 5.8% and 46.1 points on the same benchmarks. InternVL3.5-ReGATE also shows clear improvements, including +7.7/+4.5 points on MME, 0.9% on VizWiz, and 1.6% on POPE. For long-video understanding, VideoChat2-ReGATE shows strong improvements of 6.7% on Video-MME and 4.5% on MLVU. VideoLLaMA2-ReGATE also improves, though more modestly, with gains of 0.8% and 1.3% on the same tasks. At the larger scale, InternVL3.5-ReGATE further improves performance by 0.6% on Video-MME, and 0.5% on MLVU. Short video tasks benefit as well. VideoLLaMA2-ReGATE improves by 1.6% on MVBench and 1.1% on Perception, and VideoChat2-ReGATE gains 0.9% and 1.6%. InternVL3.5-ReGATE also shows improvements, with +1.3% on MVBench and +1.4% on Perception. Beyond performance metrics, ReGATE generalizes well across architectures and training strategies. It works effectively with transformer backbones such as Mistral, Qwen2, and Qwen3, and retains its benefits under both full-parameter and LoRA fine-tuning. These results confirm that ReGATE’s efficiency and accuracy gains are architecture- and training-strategy–agnostic, enabling seamless integration into diverse multimodal models.
Supplementary Material
6 Sparse Decoder Forward and Masking Details
Algorithm 1 shows how the binary mask controls computation within each decoder layer. The hidden states are first normalized, and self-attention is applied only to tokens marked active in the mask. The outputs of attention are added back through the residual connection, after which the MLP is executed solely on the active positions. Finally, the updated values are written back to the corresponding slots, while masked tokens remain unchanged and are propagated forward through residual paths. This ensures that only informative tokens are processed, while inactive ones are efficiently skipped without disrupting model integrity.
Q/K/V Masking. For clarity, we specify how the binary mask is applied within transformer layers. During self-attention, queries (Q) are computed only for the active tokens, while keys (K) and values (V) are restricted to the same set, with pruned tokens excluded from the gather step. In the feed-forward network (FFN), hidden states at active positions are passed through the MLP and the outputs are written back to their original slots. Residual connections ensure that skipped tokens simply propagate unchanged across layers. This design realizes sparse computation while leaving the model parameters intact and fully compatible with standard transformer implementations.
7 Additional Ablation Studies
In this section, we delve deeper into the design choices of ReGATE, specifically examining the sensitivity of the weighting hyperparameter and the impact of the reference teacher’s capacity.
7.1 Impact of the Weighting Factor
@@
To validate the contributions of the individual components within our dual-signal token scoring mechanism, we conduct an ablation study on the hyperparameter . This coefficient balances the two core signals in our difficulty score formulation: , where is the student’s dynamic EMA difficulty and is the static reference loss from the teacher model. We evaluate three values for : 0.0, 0.5, and 1.0. The experiments use our VideoLLaMA2-ReGATE setup with all other hyperparameters fixed for a fair comparison.
As shown in Table 7, setting achieves the best balance. Relying solely on the student’s historical difficulty () fails to capture zero-shot visual dependencies, while relying exclusively on the teacher () ignores the student’s training dynamics. The combined signal effectively isolates tokens that are both visually critical and challenging for the current model state.
@@
| Student EMA Only | 47.7 | |
| Reference Loss Only | 46.4 | |
| Combined Signals | 48.9 |
7.2 Impact of Reference Teacher Capacity
| Teacher Model | Avg. Acc. (%) |
| Qwen2-1.5B | 45.4 |
| Qwen2-57B | 46.8 |
| Qwen2-7B (Ours) | 48.9 |
A critical design choice in ReGATE is the selection of the reference teacher. We investigate how the capacity of the teacher model affects the student’s performance. Specifically, for the VideoLLaMA2-7B [cheng2024videollama2] student (based on Qwen2-7B [yang2024qwen2]), we evaluate reference signals from a smaller model (Qwen2-1.5B) and a substantially larger model (Qwen2-57B-A14B), comparing them against the matched Qwen2-7B baseline. To ensure fair comparison, we restrict the study to a single model family, which guarantees consistent tokenization and avoids misalignment between teacher and student. Cross-architecture teachers (e.g., Mistral guiding Qwen) are excluded because distinct tokenizers yield inconsistent token boundaries for the same text, violating the one-to-one mapping required for per-token reference loss computation in ReGATE. Such mismatch renders the reference signal mathematically invalid.
The results in Table 8 show that the matched Qwen2-7B teacher outperforms both the smaller 1.5B and the much larger 57B models. This pattern reveals a capacity mismatch effect: performance degrades when the teacher’s capability diverges too far from the student’s.
When using the large Qwen2-57B teacher, performance drops because the teacher is too strong relative to the student. With its vast world knowledge, the 57B model can often predict complex tokens purely from textual context, producing very low reference loss. However, the 7B student lacks this prior knowledge and instead depends on visual grounding to infer these tokens. As a result, the large teacher mistakenly signals ReGATE to prune these positions, causing the student to skip crucial computations. Conversely, the smaller Qwen2-1.5B teacher underestimates the student’s ability. It assigns high loss to linguistic patterns or factual details that the 7B student already handles easily. This floods the top- selection with linguistically difficult but visually irrelevant tokens, wasting computation on trivial cases instead of focusing on visual reasoning.
Overall, these findings suggest that the optimal reference teacher should be capacity-aligned with the student. The reference loss is most effective when it reflects the student’s own uncertainty, directing computation precisely where textual priors fall short.
@@ 7.2 Impact of Reference Teacher Capacity
| SEED | Answer with the option’s letter from the given choices directly. |
@@
| Video-MME | Answer with the option’s letter from the given choices directly. |
@@
8 Additional Benchmarks Details
Table 9 lists the evaluation prompts corresponding to each benchmark used in the experiments, most of which are adapted from LMMs-Eval [zhang2024lmmseval].
9 Qualitative Analysis of Reference Loss
To validate the core mechanism of ReGATE, we qualitatively analyze the reference loss signal that guides its token selection. We assume that a high loss score from the text-only teacher indicates that a token requires visual information to be understood. Figure 5 shows two video Q&A examples, visualizing the loss for each word in the answer as calculated by a Mistral-7B [jiang2023mistral7b] teacher model.